OpenStudy (anonymous):

Solve equation and check solution. -8(t-9)-(t+4)=1 t=

OpenStudy (anonymous):

just distribute

OpenStudy (anonymous):

t= 67/9

OpenStudy (anonymous):

how did you gt t=67/9

OpenStudy (anonymous):

the further i got was -8t-72-t-4=1

OpenStudy (anonymous):